OpenStudy (anonymous):

hey what is the gcf of 42 and 184 @skateboarding

OpenStudy (anonymous):

It is 2 use euclid's algorithm: http://codepad.org/2JVckXkQ

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I think they just want the answers and dont care how to find it agdg lol

OpenStudy (anonymous):

42=2*3*7 184=2^3*23 gcf=2

OpenStudy (anonymous):

well either euclid's algorithm or prime factorization is effective.

OpenStudy (anonymous):

Yeah i know, just thought i would save you from typing that last part lol

OpenStudy (anonymous):

over and over again